you wrote a computer program that generate two random numbers between 1 and 8. what is the probability that both values will be less than 4

Answer by stanbon(75887)

you wrote a computer program that generate two random numbers between 1 and 8. what is the probability that both values will be less than 4
------
For a single value to be less than 4:
# of ways to succeed: 3
# of random results: 8
P(a single digit is < 4) = 3/8
----
P(2 selected digits are < 4) =(3/8)^2 = 9/64
============================
Cheers,
Stan H.